2) A rock is dropped from the top of a 45m tower and at the same time, a ball is thrown from the top of the tower in a horizontal direction, Air resistance is neglected.
The ball and the rock hits the ground a distance of 30m apart.
What is the horizontal velocity of the ball thrown?
1
Expert's answer
2018-12-10T15:31:10-0500

The horizontal distance traveled by rock

L=v√(2H/g)

So, the horizontal velocity of the rock

v=L/√(2H/g)=(30 m)/√((2×45 m)/(10 m/s^2 ))=10 m/s